OpenStudy (andijo76):

Which pair has equally likely outcomes? List the letters of the two choices below which have equal probabilities of success, separated by a comma. A standard deck of cards has 12 face cards and four Aces (Aces are not face cards). A. rolling a sum of 4 on two fair six sided dice B. drawing a black seven out of a standard 52 card deck given it’s not a face card or an Ace. C. rolling a sum of 11 on two fair six sided dice D. rolling a sum of 7 on two fair six sided dice E. drawing a three out of a standard 52 card deck given it’s not a face card or an Ace.

OpenStudy (anonymous):

guess we have to find them all right?

OpenStudy (anonymous):

A: 3 out of 36 possibilities so \[\frac{3}{36}=\frac{1}{12}\]

OpenStudy (andijo76):

lol yes haha

OpenStudy (anonymous):

B is worded funny, but i guess we are supposed to think it means you know it is not a face card or an ace, so there are 12 face cards, 4 aces for a total of 16 and therefore out deck has 36 cards, 4 sevens so \[\frac{4}{36}=\frac{1}{9}\] for B

OpenStudy (anonymous):

eleven has 2 possibilities out of 36 so \[\frac{2}{36}=\frac{1}{18}\] no matches yet!

OpenStudy (anonymous):

probability you roll a seven on a pair of dice is \[\frac{1}{6}\] so it is up to the last one

OpenStudy (anonymous):

ok i made a mistake. B says "black 7" not 7 so it should be \[\frac{2}{36}=\frac{1}{18}\]

OpenStudy (anonymous):

so B and C are the same. now lets try E again 36 cards left in the deck, and 4 are threes so this time it is \[\frac{4}{36}=\frac{1}{9}\] i will let you find the matches